OpenStudy (anonymous):

A pizza is cut into 16 equal pieces. If the pizza has a diameter of 14 inches, what is the area of one piece of pizza? Round to the nearest hundredth. A. 5.5 in^2 B. 9.62 in^2 C. 14.36 in^2 D. 19.2 in^2

jhonyy9 (jhonyy9):

diameter 14 this mean the radius is 7 area = pir^2 =3,14*49 = 153,86 153,86/16 = 9,6162 so 9,62 B is right